Best Incorporated

Balance Sheet (partial)

At December 31, Year 6

Stockholders' Equity:

Preferred stock (par $100)

$100,000

Common stock (par $0.10)

52,500

Additional Paid in capital

9,136,000

Total paid in capital

9,288,500

Retained earnings

658,200

Treasury stock (3,000 common shares)

(96,000)

Total stockholders' equity

$10,042,700

Assume that the company sold 1,000 shares of its treasury stock for $35 per share.

How much would additional paid in capital change? $_______

How much would Stockholders’ Equity change?         $_______

How many shares of Common stock would be outstanding after the reissuance?    _______ shares

How many shares of Commons stock would be issued after the reissuance?     _______ shares
Answer #1

7) Cost per share of treasury stock = 96000/3000 = 32 per share

Additional paid in capital change = (35-32)*1000 = Increase by $3000

8) Stockholder's equity changes = Increase by $35000

9) Outstanding shares = (52500/.10-3000+1000) = 523000 Shares

10) Issued shares = 525000 Shares